

********************************************************************
UARS WEEKLY STATUS REPORT - 17 Apr 1998
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM was turned off on April 15 and should resume operations on
April 20. Otherwise, SUSIM successfully gathered its normal solar,
offset, and occultation scan data this week without apparent problems.
***********
PEM
***********
PEM is currently off.
PEM is awaiting evaluation by GSFC of its proposed activation procedure.
The minimal configuration for PEM AUX activation is the zenith VMAG, MEPS1,
MEPS2, and HEPS1 units, only. This configuration is driven by the capabilities
of the spacecraft OBC and may increase if OBC program area increases. PEM
activation is proposed to be OBC driven when there exists plenty of power to
operate PEM. It is anticipated that the OBC will activate PEM on the daylight
portions of every orbit.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
MLS was reactivated successfully on Apr 15, in the reduced 85 watt power
mode. MLS instrument operations have continued flawlessly. MLS is scheduled to
operate through Apr 19.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
The following paper was just accepted for publication in JGR, and the
camera-ready copy submitted.
Tidal influence on mid-latitude airglow: Comparison of satellite and
ground-based observations with TIME-GCM predictions
G.G. Shepherd, R.G. Roble, S.-P. Zhang, C. McLandress and R.H. Wiens
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF will be unavailable during the scheduled power outage in building 23
occurring the afternoon of April 24 to the morning of April 27. Level 0 files
which will not be received during this time period will have to be rescheduled
from White Sands.
The CDHF has received a total of 381 CD-ROM requests as of 17-April-1998.
There were no new orders received.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
25 days PEM (3.32) PROD (1379R-1124C=255L)
9 days CLAES (7.0) ENG (72R-49C-1F=22L)
764 days ATT (7.2) PROD (2178R-1203C=975L)
2 days HAL (4.80) ENG (7R-7C=0L)
The CDHF continued work on the CDHF Year 2000 compliance matrix and found out
the Y2K patches for VMS are out for VAX 5.5-2 and Alpha 6.1.
CDHF pending software upgrades: IDL 5.0, OpenIngres, UCSS.
***********
FOT
***********
REPORTING PERIOD
This report covers 03 April 1998 (Orbit 35854, GMT Day 093) through 10
April 1998 (Orbit 35959, GMT Day 100).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
WINDII 094/0029 35853 On
HRDI 095/0033 35869 On
WINDII 097/2349 35914 Off
HALOE 098/0235 35916 On
A summary of Instrument operations supplied by the Mission Planning Group
for the month of March 1998 is contained in Appendix F, UARS Instrument
Operations Summary.
The beta angle is now increasing toward a maximum angle of 45.7 degrees on
19 April 1998 (DOY 109).
Two manually generated SSPP Stellar Offset experiments requested by
SOLSTICE were performed on GMT days 097 and 099.
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 24.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 1.1 degrees C. Additional battery performance data for
each battery on 04 April 1998 and 09 April 1998 are provided in Appendix D.
The clock error ranged from -10.4 to +9.8 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = none
% OUT-OF-LIMITS = 07
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 02
% HRDI 01
% HALOE 01
% WINDII 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 01
One (1) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/062 - 2287.496945
Transponder B: 98/062 - 2287.497010
GROUND SYSTEM OPERATION
One (1) operational support problem(s) occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/099 show the observatory data loss to be
20 hours, 42 minutes, 36.532 seconds (an increase of 3.392 seconds since
last report period). This is a 0.0361 percent data loss which equals a
99.9639 percent data capture for the mission.
The increase of 3.392 seconds occurred over a two day period as follows:
7 2.336 seconds on Day 095 due to insufficient overlap, and
7 1.056 seconds on Day 098 due to hits in the data.
SOLSTICE operations has submitted a requested performance of a stellar
offset experiment on GMT Day 104.
The FOT continues to experience ephemeris transition comparisons in excess
of 1.5 km when processing the ephemeris in the Command Management System
(CMS) software. The FOT was able to get a response from FDF regarding the
possible need to transition to more frequent spacecraft ephemeris
deliveries (i.e. back to a daily delivery schedule instead of the current
Monday/Wednesday/Friday schedule). The FOT spoke to Craig Woodruff (who is
now only working one day per week as an FDF analyst) who requested the FOT
to fax to FDF several memorandum and supporting documentation outlining the
ephemeris delivery requirements. The FDF analyst will review the documents
and make a decision as to if and when the increased ephemeris delivery
transition will be necessary.
No IP testing was performed by the UARS FOT this week. However, the FOT
met with Tom Marsh to discuss the details of configuring a forward link
command loop-back test using the UARS IP conversion devices (CDs) located
in the MSOCC area. Apparently the previously discussed loop-back test
(using RFSOC) may be too difficult to perform. The suggested test would
involve the FOT sending commands from the AP and having the commands routed
through the CD, where they would be converted to IP format. Then, the
commands would be looped back through the CD (where they would be converted
back to 4800-bit blocks) and sent to the UTTS for command verification.
While Mr. Marsh works out the remaining details of this test, the FOT plans
to continue testing the return link IP path, concentrating particularly on
Q-channel OBC dumps.
The FOT continued making modifications to the OBCSTF resident POCC
database. These changes will be incorporated into the next operational
database release, version 66.
OTHER SIGNIFICANT EVENTS
Work on the upgrade of the OBC Power Monitor (PMON) software continued.
FOT continues to support CNMOS in their efforts to make necessary software
changes (MSOCC, CMS, DCF, etc.) to support the year 2000 rollover.
Monitoring of the magnetic field model performance continues.
FOT continues to improve a Power Profile for UARS based on the current
status of the Modular Power System.
FOT continues to support the new UARS ground system re-engineering effort.
MEETINGS
No UARS Battery telecon was held this week. The next telecon has not been
scheduled.
FOT ( Erik Larson) supported the Sixteenth Annual Space Power Workshop in
Long Beach, CA. FOT presented RUARS Program ExperienceS briefing,
describing in-flight experience with the UARS power subsystem, and
recommendations for future spacecraft design based upon lessons learned
with UARS.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is currently scheduled for 05 May 1998.
A second building 23 power outage is scheduled from about 12 p.m., 24 April
1998, until about 12 p.m., 27 April 1998.
The UARS orbit altitude is dropping slowly and a drag makeup orbit adjust
is targeted for 05 August 1998.
PERSONNEL CHANGES
None.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
Appendix F, UARS Instrument Operations Summary
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 10 April 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
35854 094/0029 98-123 WINDII Power ON
35869 095/0033 98-124 HRDI Power ON
35914 097/2349 98-124 WINDII Power OFF
35916 098/0236 98-126 HALOE Power ON
35920 098/1015 98-127 OBC Clk Adj fr 0.00 to 20.6 msec/day
35940 099/1819 98-128 OBC Clk Adj fr 20.6 to 0.00 msec/day
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 10 April 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPSEQBIMON Y-HI HRDI QB Current Monitor 35818,59
HRPTSPBIMON Y-HI HRDI Tele PB Current Monitor 35871,72,88,903,18,24
NBRECBI Y-HI NBTR B Record Current 35859
OBSSOC2 Y-LO Battery 2 Current SOC 35918
QTSISAIF Y-HI SSPP Baseplate by SISA I/F 35843-46, 50-53,58-68,
74,80,89,90-98,903,
09,918,19, 43
WIEUTEMP Y-LO WINDII EU Temperature Continuous
WIQBCURRENT Y-LO WINDII Quiet Bus Current 35854,56,60-69
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 10 April 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
35920 098/10:20 (08 Apr 98) from 00.0 to 20.6 msec/day
35940 099/18:26 (09 Apr 98) from 20.6 to 00.0 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 10 April 1998
04 Apr 1998 - GMT Day 094 Beta = 8.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/ +0.0 18.4 0.000 0.0
2 +112.0/ +0.0 +6.44/+4.83 +32.0/-15.2 24.0 1.023 22.9
3 +78.4/ -5.6 +5.15/+3.87 +31.2/-16.0 24.0 1.021 22.4
09 Apr 1998 - GMT Day 099 Beta = 26.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-3.87 +0.0/ +0.0 18.4 0.000 0.0
2 +61.6/ -11.2 +6.80/+5.30 +31.8/-15.4 24.8 1.021 23.0
3 +67.2/ -5.6 +5.47/+4.51 +31.8/-16.1 24.8 1.024 22.5
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes
APPENDIX E
UARS ANOMALY REPORTS
Week ending 10 April 1998
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-016 FOT 35882 095/2020 14170 TAC Crashed
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-016 FOT 35882 095/20:20 TAC 4 Crash
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-015 Generic Late Acq for April 1998 (Open until 30 April 1998).
_______________________________________________________________
APPENDIX F
UARS INSTRUMENT OPERATIONS SUMMARY
Week ending 10 April 1998
INSTRUMENT SUMMARY FOR MONTH OF March 1998 (hours)
UNPLANNED UNSCHEDULED
ON OFF OFF OFF
(Insufficient (Other than
Instrument (Actual) (Scheduled) Power) Power Problem) COMMENTS
---------- ------ ----------- ----------- ------------- ------------
ACRIM 622 122 0 0
HALOE 524 220 0 0
HRDI 380 364 0 0
MLS 310 434 0 0
PEM (axis) 117 627 0 0
PEM (zeps) 0 744 0 0
PEM (neps) 0 744 0 0
SOLSTICE 744 note 1 0 0 Note 1
SUSIM 607 137 0 0
WINDII 355 389 0 0
Note 1 - No star data March 12-16
Cold side of spacecraft was facing northward: 0.0 days
Cold side of spacecraft was facing southward: 31.0 days